# CodeWiki: Automated Repository-Level Documentation at Scale

### v1

@@ Abstract
Developers spend nearly 58% of their time understanding codebases, yet maintaining comprehensive documentation remains challenging due to complexity and manual effort. While recent Large Language Models (LLMs) show promise for function-level documentation, they fail at the repository level, where capturing architectural patterns and cross-module interactions is essential. We introduce CodeWiki, the first open-source framework for holistic repository-level documentation across seven programming languages. CodeWiki employs three innovations: (i) hierarchical decomposition that preserves architectural context, (ii) recursive agentic processing with dynamic delegation, and (iii) synthesis of textual and visual artifacts including architecture diagrams and data flows. We also present CodeWikiBench, the first repository-level documentation benchmark with multi-level rubrics and agentic assessment. CodeWiki achieves 68.79% quality score with proprietary models and 64.80% with open-source alternatives, outperforming existing closed-source systems and demonstrating scalable, accurate documentation for real-world repositories.
@@ 1 Introduction
In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity. Studies indicate that developers spend approximately 58% of their working time on program comprehension activities (Xia et al., 2018), highlighting the critical importance of accessible and accurate documentation. Despite documentation being recognized as essential for software maintenance and collaboration (de Souza et al., 2005; Zhi et al., 2015; Chen and Huang, 2009), the manual creation and maintenance of repository-level documentation remain a labor-intensive and often neglected aspect of software engineering practice (McBurney et al., 2018). The emergence of Large Language Models (LLMs)  (Zhang et al., 2023; He et al., 2025b) has opened opportunities for automating documentation generation tasks. However, current approaches still face two fundamental limitations.
@@
Recent advances in code understanding and natural language generation have demonstrated promising results in function-level and file-level documentation tasks (Feng et al., 2020; Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025) However, existing approaches face significant challenges when scaling to repository-level documentation.
Unlike function-level documentation that focuses on individual components, repository-level documentation task is to capture architectural patterns, cross-module interactions, data flows, and system-wide design decisions (Rai et al., 2022; Treude et al., 2020). Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance (Nassif and Robillard, 2025).
@@
Despite progress, systematic benchmarks remain lacking. Function-level and file-level work relies on general-purpose metrics (BLEU, ROUGE) that fail to capture documentation quality nuances (Roy et al., 2021; Evtikhiev et al., 2023). Repository-level evaluation is particularly challenging as documentation admits multiple valid structures sharing underlying architectural content, highlighting the need for robust assessment beyond surface-level similarity.
@@ 2.1 Repository Analysis and Hierarchical Module Decomposition
Following DocAgent (Yang et al., 2025), we employ Tree-Sitter parsers to extract ASTs, systematically identifying functions, methods, classes, structs, modules, and interdependencies: function calls, class inheritance, attribute access, and module imports. We normalize these to a unified depends_on relation for cross-language generalization, creating directed graph $G=(V,E)$ where edge $(u,v)\in E$ indicates component $u$ depends on $v$ (Liu et al., 2023).
@@ 2.2 Recursive Documentation Generation
Initialize documentation workspace;
for each leaf module $m$ in $T$ do
   agent $\leftarrow$ CreateAgent($m$, $G$, $T$);
 while $m$ not fully processed do
      agent.GenerateDoc($m$.name);
    if delegation requested then
         sub $\leftarrow$ agent.GetDelegationSpecs();
         UpdateModuleTree($T$, $m$, sub);
       foreach new leaf $m^{\prime}$ in sub do
            Recursively process $m^{\prime}$;
        ReviseDoc($m$.name);
ProcessParentModules($T$);
return Repository documentation;
@@ 3 The First Benchmarking System - CodeWikiBench
Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization. Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels. Recent research highlights traditional metrics’ limitations for code tasks (Evtikhiev et al., 2023; Haldar and Hockenmaier, 2024), while LLM-as-a-Judge paradigms offer new possibilities (Wang et al., 2025; Crupi et al., 2025; He et al., 2025a). We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
3.1 Hierarchical Rubric Generation
Following (Starace et al., 2025), we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2. We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax. This captures multi-level technical documentation from architectural overviews to implementation details.
@@ 4.2 Baseline
To the best of our knowledge, CodeWiki represents the first open-source framework specifically designed for comprehensive repository-level documentation generation. The absence of comparable open-source solutions necessitates comparison with closed-source alternatives. We select DeepWiki  (DeepWiki, 2025) as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
@@ 4.3 Experimental Setup
Repository Selection
We carefully curate a comprehensive benchmark consisting of 21 open-source repositories that represent diverse programming languages, project scales, and application domains. The selection criteria prioritize: (1) language diversity to test cross-language generalization, (2) varying codebase sizes from 16K to 1.7M lines of code, (3) active maintenance with comprehensive official documentation for rubric generation, and (4) diverse application domains including web frameworks, automation tools, machine learning platforms, and system utilities.
@@
Model Configuration
We evaluate CodeWiki with Claude Sonnet 4 (CodeWiki-sonnet-4) and Kimi K2 Instruct (CodeWiki-kimi-k2). Rubric generation uses multiple models (Claude Sonnet 4, Gemini 2.5 Pro, Kimi K2 Instruct) to reduce bias. Evaluation employs three judge models (Gemini 2.5 Flash, GPT OSS 120B, Kimi K2 Instruct) for independent assessments (details in Appendix E).
Evaluation Protocol
For each repository, we execute the complete CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides. The evaluation process follows our hierarchical rubric methodology: (1) generating repository-specific rubrics from official documentation, (2) performing agentic assessment of generated documentation, and (3) computing weighted scores with uncertainty quantification through standard deviation tracking.
@@ 4.4.2 RQ2: Cross-Language Generalization
Our analysis reveals distinct performance patterns across programming language categories, as visualized in Figure 3. CodeWiki-sonnet-4 demonstrates exceptional performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%. CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
@@ B.1.1 Existing Approaches
Advances in natural language processing (NLP) have enabled significant progress in the task of code summarization. Existing methods  (Khan and Uddin, 2023; Poudel et al., 2024; Makharev and Ivanov, 2025; Lomshakov et al., 2024; Ahmed and Devanbu, 2023; Choi et al., 2023; Luo et al., 2024; Yang et al., 2025) primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
@@
Early neural approaches established the foundation for automated documentation. CodeBERT (Feng et al., 2020) demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements. For instance, (Khan and Uddin, 2023) reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods. DocuMint (Poudel et al., 2024) further explored the use of small language models for Python docstring generation with promising results. However, these approaches primarily rely on local code context and do not leverage repository-level information to improve documentation quality.
@@
Recognizing the limitations of function-only context, recent work has explored broader codebase information to enrich documentation.  (Makharev and Ivanov, 2025) demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation. ProConSuL (Lomshakov et al., 2024) incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations. RepoAgent (Luo et al., 2024) introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth. DocAgent (Yang et al., 2025) advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
@@ B.1.2 Existing Evaluation Methods
$\hookrightarrow$ Limitations and Our Motivation: While these methods provide a lightweight quantitative measure, they often fail to capture deeper semantic aspects of the documentation, such as clarity, correctness, and completeness in explaining the underlying code behavior. Recent studies (Roy et al., 2021; Evtikhiev et al., 2023) have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality. This gap highlights the need for an assessment system that can fairly and accurately evaluate documentation, ensuring that the generated content truly supports developers in comprehending the system.
@@ B.2 Rubric-based evaluation
A repository can be described through multiple representations, each using different structures and presentation styles while sharing the same underlying architecture. This diversity makes it difficult to evaluate open-ended outputs automatically, as no single metric can reliably evaluate which version is better. Although human evaluation provides trustworthy results, it is resource-intensive and slow. To overcome these limitations and facilitate rapid assessment, many benchmarks (Starace et al., 2025; Lin et al., 2025; Sirdeshmukh et al., 2025; Arora et al., 2025) often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs. In our CodeWikiBench benchmark, we similarly construct a set of hierarchical rubrics based on the official documentation, providing a structured and fine-grained framework to evaluate the quality and completeness of generated repository-level documentation.
@@ C.3 Overall Reliability Score
Repository Semantic (%) Structural (%) All-Hands-AI–OpenHands 73.64 68.82 sveltejs–svelte 75.56 68.88 puppeteer–puppeteer 72.97 76.36 Unity-Technologies–ml-agents 75.95 71.22 elastic–logstash 73.13 68.22 wazuh–wazuh 72.20 67.62 electron–electron 72.12 74.79 Average 73.65 70.84

### v2

# CodeWiki: Evaluating AI’s Ability to Generate Holistic Documentation for Large-Scale Codebases

@@ Abstract
Given a large and evolving codebase, the ability to automatically generate a holistic, architecture-aware documentation that captures not only individual functions but also their cross-file, cross-module, and system-level interactions remains an open challenge. Comprehensive documentation is essential for long-term software maintenance and team collaboration, yet current automated approaches still fail to model the rich semantic dependencies and architectural structures that define real-world software systems. We present CodeWiki, a unified framework for automated repository-level documentation across seven programming languages. CodeWiki introduces three key innovations: (i) hierarchical decomposition that preserves architectural context across multiple levels of granularity, (ii) recursive multi-agent processing with dynamic task delegation for scalable generation, and (iii) multi-modal synthesis that integrates textual descriptions with visual artifacts such as architecture diagrams and data-flow representations. To enable rigorous evaluation, we introduce CodeWikiBench, a comprehensive benchmark featuring multi-dimensional rubrics and LLM-based assessment protocols. Experimental results demonstrate that CodeWiki achieves a 68.79% quality score with proprietary models, outperforming the closed-source DeepWiki baseline (64.06%) by 4.73%, with particularly strong improvements on high-level scripting languages (+10.47%). We open-source CodeWiki to foster future research and community adoption.
[…]
@@ 1 Introduction
In the rapidly evolving landscape of software development, maintaining comprehensive and up-to-date documentation has become increasingly challenging as codebases grow in size and complexity. Recent industry surveys reveal that documentation represents a critical bottleneck in modern development workflows: approximately 31% of developers currently rely heavily on AI for documenting code, while 25% use AI extensively for creating and maintaining documentation—making documentation tasks among the highest priorities for AI integration over the next 3-5 years [25]. Notably, while developers express high satisfaction with AI tools for searching and learning, they indicate plans to substantially increase AI adoption for documentation and testing tasks, suggesting current tools remain insufficient for these critical needs. This widespread focus on AI-driven documentation reflects its dual burden: while essential for software maintenance and collaboration [6, 31, 3], manual creation and maintenance remain labor-intensive and often neglected [19]. The emergence of Large Language Models (LLMs) [30, 12] has opened opportunities for automating documentation generation. However, current approaches still face two fundamental limitations.
@@
Recent advances in code understanding and natural language generation have demonstrated promising results in function-level and file-level documentation tasks [9, 13, 21, 18, 16, 1, 4, 17, 29] However, existing approaches face significant challenges when scaling to repository-level documentation.
Unlike function-level documentation that focuses on individual components, repository-level documentation task is to capture architectural patterns, cross-module interactions, data flows, and system-wide design decisions [22, 27]. Furthermore, the hierarchical nature of software systems requires documentation that can serve multiple audiences, from high-level stakeholders seeking architectural overviews to developers requiring detailed implementation guidance [20].
@@
Despite progress, systematic benchmarks remain lacking. Function-level and file-level work relies on general-purpose metrics (BLEU, ROUGE) that fail to capture documentation quality nuances [23, 8]. Repository-level evaluation is particularly challenging as documentation admits multiple valid structures sharing underlying architectural content, highlighting the need for robust assessment beyond surface-level similarity.
@@ 2.1 Repository Analysis and Hierarchical Module Decomposition
Following DocAgent [29], we employ Tree-Sitter parsers to extract ASTs, systematically identifying functions, methods, classes, structs, modules, and interdependencies: function calls, class inheritance, attribute access, and module imports. We normalize these to a unified depends_on relation for cross-language generalization, creating directed graph $G=(V,E)$ where edge $(u,v)\in E$ indicates component $u$ depends on $v$ [15].
@@ 2.2 Recursive Documentation Generation
1 Initialize documentation workspace;
2 for each leaf module $m$ in $T$ do
3   agent $\leftarrow$ CreateAgent($m$, $G$, $T$);
4 while $m$ not fully processed do
5      agent.GenerateDoc($m$.name);
6    if delegation requested then
7         sub $\leftarrow$ agent.GetDelegationSpecs();
8         UpdateModuleTree($T$, $m$, sub);
9       foreach new leaf $m^{\prime}$ in sub do
10            Recursively process $m^{\prime}$;
11
12         end foreach
13        ReviseDoc($m$.name);
14
15      end if
16
17   end while
18
19 end for
20ProcessParentModules($T$);
21 return Repository documentation;
@@ 3 The First Benchmarking System - CodeWikiBench
Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization. Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels. Recent research highlights traditional metrics’ limitations for code tasks [8, 10], while LLM-as-a-Judge paradigms offer new possibilities [28, 5, 11]. We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
3.1 Hierarchical Rubric Generation
@@
Following [26], we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2. We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax. This captures multi-level technical documentation from architectural overviews to implementation details.
@@ 4.2 Baseline
To the best of our knowledge, CodeWiki represents the first open-source framework specifically designed for comprehensive repository-level documentation generation. The absence of comparable open-source solutions necessitates comparison with closed-source alternatives. We select DeepWiki [7] as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
@@ 4.3 Experimental Setup
4.3.1 Repository Selection
We carefully curate a comprehensive benchmark consisting of 21 open-source repositories that represent diverse programming languages, project scales, and application domains. The selection criteria prioritize:
  1. 1.
    Language diversity to test cross-language generalization
  2. 2.
    Varying codebase sizes from 16K to 1.7M lines of code
  3. 3.
    Active maintenance with comprehensive official documentation for rubric generation
  4. 4.
    Diverse application domains including web frameworks, automation tools, machine learning platforms, and system utilities
@@ 4.3.1 Repository Selection
4.3.2 Model Configuration
We evaluate CodeWiki with two different LLM backends:
  • CodeWiki-sonnet-4: Uses Claude Sonnet 4 (proprietary model)
  • CodeWiki-kimi-k2: Uses Kimi K2 Instruct (open-source model)
For rubric generation, we employ multiple models (Claude Sonnet 4, Gemini 2.5 Pro, Kimi K2 Instruct) to reduce single-model bias. The evaluation process uses three independent judge models (Gemini 2.5 Flash, GPT OSS 120B, Kimi K2 Instruct) for robust assessment. Additional implementation details are provided in Appendix E.
4.3.3 Evaluation Protocol
For each repository, we execute the complete CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides. The evaluation process follows our hierarchical rubric methodology:
  1. 1.
    Rubric Generation: Repository-specific rubrics derived from official documentation
  2. 2.
    Agentic Assessment: Multiple judge agents evaluate generated documentation against rubric criteria
  3. 3.
    Score Aggregation: Weighted scores computed with uncertainty quantification through standard deviation tracking
@@ 4.4.2 RQ2: Cross-Language Generalization
Our analysis reveals distinct performance patterns across programming language categories, as visualized in Figure 3.
@@
CodeWiki-sonnet-4 demonstrates exceptional performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%. CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
@@ B.1.1 Existing Approaches
Advances in natural language processing (NLP) have enabled significant progress in the task of code summarization. Existing methods [13, 21, 18, 16, 1, 4, 17, 29] primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
@@
Early neural approaches established the foundation for automated documentation. CodeBERT [9] demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements. For instance, [13] reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods. DocuMint [21] further explored the use of small language models for Python docstring generation with promising results. However, these approaches primarily rely on local code context and do not leverage repository-level information to improve documentation quality.
@@
Recognizing the limitations of function-only context, recent work has explored broader codebase information to enrich documentation. [18] demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation. ProConSuL [16] incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations. RepoAgent [17] introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth. DocAgent [29] advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
@@ B.1.2 Existing Evaluation Methods
$\hookrightarrow$ Limitations and Our Motivation: While these methods provide a lightweight quantitative measure, they often fail to capture deeper semantic aspects of the documentation, such as clarity, correctness, and completeness in explaining the underlying code behavior. Recent studies [23, 8] have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality. This gap highlights the need for an assessment system that can fairly and accurately evaluate documentation, ensuring that the generated content truly supports developers in comprehending the system.
@@ B.2 Rubric-based evaluation
A repository can be described through multiple representations, each using different structures and presentation styles while sharing the same underlying architecture. This diversity makes it difficult to evaluate open-ended outputs automatically, as no single metric can reliably evaluate which version is better. Although human evaluation provides trustworthy results, it is resource-intensive and slow. To overcome these limitations and facilitate rapid assessment, many benchmarks [26, 14, 24, 2] often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs. In our CodeWikiBench benchmark, we similarly construct a set of hierarchical rubrics based on the official documentation, providing a structured and fine-grained framework to evaluate the quality and completeness of generated repository-level documentation.
@@ C.3 Overall Reliability Score
Repository Semantic (%) Structural (%)
All-Hands-AI–OpenHands 73.64 68.82
sveltejs–svelte 75.56 68.88
puppeteer–puppeteer 72.97 76.36
Unity-Technologies–ml-agents 75.95 71.22
elastic–logstash 73.13 68.22
wazuh–wazuh 72.20 67.62
electron–electron 72.12 74.79
Average 73.65 70.84